Grade 12Mechanics

A body falls from a height h=200m.The ratio of distance travelled in each 2 s, during t=0 to t=6s of the journey is

Vaibhav gaikwad
9 Years agoGrade 12

1 Answer

Khimraj
9 Years ago
Distance in first two seconds is (1/2)g(2)² = 20 in first four seconds is (1/2)g(4)² = 80 and in first six seconds is (1/2)g(6)² = 180 so ratio of distance traveled in each two seconds is 20:(80-20):(180-80) = 20:60:100 = 1:3:5.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer